Corrections to “A survey of random random walks on finite groups” Probab. Surveys 2 (2005), 33-63

On p. 44, in the display just below the line with “Let’s deal with the case k = 2 now”, P̂_{a_{1},a_{2}}(j) should be |P̂_{a_{1},a_{2}}(j)|^{2}.

Thanks to Joseph McCollum for his help in finding this point.